OpenStudy (anonymous):

Find an equation for the circle of curvature of the curve r(t) = (2 ln(t))i - [t-1/t]j at the point (0, -2) , where t = 1

OpenStudy (ybarrap):

Is $$ \large{ r(t)=\left (2\ln t,-t+\cfrac{1}{t}\right )\\ } $$ ? Because $$ r(1)=(0,0) $$ and $$ r(1)\ne(0,-2) $$ Which is what you have above. You'll need $$ \large{ r'(t)=(x'(t),y'(t))\\ r''(t)=(x''(t),y''(t)) } $$ To get $$ \large{ \kappa(t)= \left|\frac{x'y''-y'x''}{({x'^2+y'^2})^{3/2}}\right|\\ } $$ http://en.wikipedia.org/wiki/Curvature#Example